OpenStudy (anonymous):

1.A mysterious box is delivered to the dinner party you are attending. The label on the box says that the volume of a box is the function f(x) = x^3 + 3x^2 – 10x – 24. To open the box, you need to identify the correct factors of f(x). Party-goers offer up solutions, and it is your job to find the right ones. Their suggestions are: •(x – 1) •(x + 2) •(x – 3) •(x + 4) •(x + 6) •(x – 12) List the correct factors. Then justify your selections with complete sentences.

OpenStudy (anonymous):

@austinL

OpenStudy (anonymous):

You can use synthetic division to do this

OpenStudy (anonymous):

although i don't even know how to do the box in the left on word for it.. does anyone know

OpenStudy (campbell_st):

just use the factor theorem which means, substitute a possible zero into the polynomial, when simplified, if the answer is zero then you have a zero and hence factor of the polynomial. as an example look at the possible factor (x + 2) which means x = -2 is a possible zero. then you need to find f(-2) and see if its zero. so \[f(-2) = (-2)^3 + 3\times (-2)^2 -10 \times(-2) - 24 = -8 + 12 + 20 -24\] and when you evalute it you find f(-2) = 0 so then you can say (x + 2) is a factor of the polynomial. here is a link to a site that explains the factor theorem hope this helps

OpenStudy (anonymous):

ohh, but how come when I just tried it with synthetic division, it wasn't a factor..?

OpenStudy (anonymous):

for (x-2)

OpenStudy (anonymous):

and btw u didn't put a link^^

OpenStudy (campbell_st):

well I showed (x + 2) is a factor.... not (x -2) and I showed it was a factor by substituting x = -2 into the cubic. here is another clue look at the cubic f(x) = (x +2)(x +a)(a + b) the product of the 3 roots and the constant term in the cubic \[2 \times a \times b = -24\] so you are looking for a pair of factors where the constants multiply to -12.. so it will eliminate a few pairs.

OpenStudy (anonymous):

it was a typo

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

so which factor do u think i should test next

OpenStudy (campbell_st):

well I would say not (x -1) so I'd say (x-3) and then x + 4) perhaps

OpenStudy (campbell_st):

so substitute f(3) then f(-4) and see if they equal zero

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

help.. can you explain what everything in the formula means f(x) = (x +2)(x +a)(a + b) the a, b, and x

OpenStudy (anonymous):

im doing (x+3) now. So I know -3 will be x, but what's a and b?

OpenStudy (anonymous):

oh wait.. i just plug it in to the original equation right

OpenStudy (anonymous):

Idk if I did it right.. but i just tried it and i got -1224 ...

OpenStudy (campbell_st):

well you know (x + 2) is a factor now text (x -3) by substituting x = 3 into the equation if its equal to zero you'll have another factor. then f(x) = (x +2)(x -3) and the last one should be (x + 4) to explain say something like, by using the factor theorem, which requires the possible zeros of the cubic to be substituted, I found f(-2) = 0 which meant (x + 2) is a factor.... then tested f(3) and found it was equal to zero so (x -3) was a factor so f(x) = (x +2)(x -3)(x + a) the product of 2, -3, and a has to equal -24 so set up an equation -6a = -24 and solve for a a = 4 so the factors are f(x) = (x +2)(x -3)(x + 4) hope this helps

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

where did the " f(x) = (x +2)(x -3)" come from?

OpenStudy (anonymous):

wait.. r u sure this is called the factor theorem? cuz this doesn't look like it: http://www.purplemath.com/modules/factrthm.htm

OpenStudy (campbell_st):

well is x +2 is a factor then x = -2 is a zero and (x -3) is a factor then x = 3 is a zero f(-2) = 0 is shown above so look at f(-3) \[f(-3) = (3)^3 + 3(3)^2 -10(3) - 24 = 27 + 27 - 30 - 24 = 0\] so x = 3 is a zero and hence (x - 3) is a factor so now you have 2 factors f(x) = (x +2)(x -3) just find the last

OpenStudy (anonymous):

i keep doing it wrong

OpenStudy (campbell_st):

well if you read the 2 examples in the link...purplemath they used synthetic division to so.... x = -4 is a zero and hence x + 4 is a factor in example 2. in example 1 the are asking to test if x - 1 is a factor, so they substituted the zero x = 1 and found it wasn't a factor... then supported the claim with synthetic division. and it is the same...

OpenStudy (anonymous):

ok i think i get it now thanks i gtg

OpenStudy (campbell_st):

so if you look at the quadratic \[f(x) = x^2 - x - 6\] it can be factored to \[f(x) = (x -3)(x + 2)\] the equation has been rewriten in factored form so to find the values of x that make the equation equal zero you need to solve x - 3 = 0 therefore x = 3 x + 2 = 0 therefore x = -2 either value will make the equation equal zero..